# Hierarchical self-assembly for high-yield addressable complexity at fixed conditions

### v1

@@ Abstract
There is evidence that the self-assembly of complex molecular systems often proceeds hierarchically, by first building subunits that later assemble in larger entities, in a process that can repeat multiple times. Yet, our understanding of this phenomenon and its performance is limited. Here we introduce a simple model for hierarchical addressable self-assembly, where interactions between particles can be optimised to maximise the fraction of a well-formed target structure, or yield. We find that a hierarchical strategy leads to an impressive yield up to at least five generations of the hierarchy, and does not require a cycle of temperatures as used in previous methods. High yield is obtained when the microscopic interaction decreases with the scale of units considered, such that the total interaction between intermediate structures remains identical at all scales. We provide thermodynamic and dynamical arguments constraining the interaction strengths where this strategy is effective. Overall, our work characterizes an alternative strategy for addressable self-assembly at a fixed temperature, and provides insight into the mechanisms sustaining hierarchical assembly in biology.
@@ I Introduction
The ribosome, the central machinery that translates messenger RNA into proteins, illustrates the surprising ability of multi-components systems to self-assemble. It is a protein complex made of $\sim 100$ proteins (in addition to ribosomal RNA), containing around $10^{4}$ amino acids. Considering the astronomical number of possible configurations these amino acids could take, how does this system self-assemble correctly? Clearly, this process is simplified by the fact that amino-acids first fold into proteins, which later form a complex. There is evidence that such a hierarchical strategy, where small units first form before assembling into larger ones, is at play in the folding of individual proteins itself Dill et al. (2008). In this view, proteins first nucleate secondary structures such as beta sheets or alpha helices Baldwin and Rose (1999) or more complex ”foldons” Fersht (1997) which later organize into the full tertiary structure. Hierarchical self-assembly is also used in chemistry to obtain supra-molecular structures, typically held together with hydrogen bonds or Van der Waals interactions, much weaker than the covalent bonds organizing molecules Elemans et al. (2003). More generally, a hierarchical organization is central to a variety of complex systems and processes, from engineering to the composition of companies from smaller entities Simon (1962). Despite its importance and ubiquity, understanding the principles governing hierarchical assembly remains a challenge.
@@
Such principles could be used and tested in the context of “addressable” self-assembly, where each unit of a target structure is distinct and must assemble into a particular location Rothemund and Andersen (2012); Rothemund (2006); Sacanna et al. (2013); Wang et al. (2012); Rogers and Manoharan (2015); Rogers et al. (2016); Hong et al. (2017); Murugan et al. (2015). Using DNA strands as building blocks with highly specific interactions, experimentalists have so realized highly complex structures containing up to tens of thousands of units Wei et al. (2012); Ke et al. (2012); Ong et al. (2017); Evans et al. (2024). A challenge is to find the particular building blocks King et al. (2024) or protocols Goodrich et al. (2021); Bupathy et al. (2022); Engel et al. (2023) such that the system assembles into a desired structure. Experiments operate in the regime where all pair interactions have roughly the same strength Hormoz and Brenner (2011), although the importance of displaying a range of interactions has been proposed in various contexts Jhaveri et al. (2024); Kaplan et al. (2014). When these interactions are strong, or equivalently the temperature is low, the system assembles many partially-formed fragments of the target structures. These fragments must break apart to assemble into copies of the target structure, but they cannot do so on accessible timescales – although the target structure has the lowest free energy, it is not kinetically accessible. In contrast, with weaker interactions the target state no longer has the lowest free energy. One solution, identified experimentally and later rationalized theoretically Reinhardt and Frenkel (2014); Jacobs et al. (2015) is to anneal temperature: start at a high temperature where nucleation is rare, then slowly lower temperature to grow the nuclei, repeating until one forms a target structure. While this protocol has worked well as a method to form one or a small number of copies of a target structure, it wastes monomers, since many monomers do not assemble into a target structure, and furthermore it requires a precise temperature protocol.
@@
$1234$
Figure 1: Schematic of our model: particles on a square lattice interacting with pairwise energies. They move using the VMMC algorithm, which moves clusters as a unit with appropriate diffusivities. This schematic contains enough monomers to make four 4-squares. Only one 4-square has fully formed.
[…]
Figure 2: Building large squares with addressable complexity using hierarchical interactions. Square monomers with sticky sides assemble into 4-squares with strong edge interactions $\delta_{1}$, 4-squares assemble into 16-squares with weaker individual edge interactions $\delta_{2}$, 16-squares assemble into 64-squares using even weaker edge interactions $\delta_{3}$, and so on. We expect good assembly when the interactions are chosen hierarchically $\delta_{2}{=}\delta_{1}/2$, $\delta_{3}{=}\delta_{2}/2$, $\delta_{4}{=}\delta_{3}/2$, etc (for appropriate values of $\delta_{1}$, as studied below). This way, the energy for gluing together $4^{k}$-squares into $4^{k+1}$-squares, is the same as for gluing together monomers into 4-squares.
@@ II.1 Model
We consider a system of sticky squares (monomers) on a two-dimensional square lattice with periodic boundary conditions, as depicted in Fig. 1. We assume that monomer $i$ interacts with monomer $j$ with an interaction energy 00 if the squares are not touching, $\infty$ if the squares overlap, and $-\delta_{ij}^{\vec{l}}$ if the squares are exactly touching with separation $\vec{l}\in\{(1,0),(-1,0),(0,1),(0,-1)\}$. The target structure is an $n$-square containing $n=2^{k}\times 2^{k}=4^{k}$ monomers (Figure 2) where $k$ is some integer, with each monomer in a distinct location. Our goal is to make $N_{c}$ copies of this target structure at a constant temperature $T$. We measure interaction energies in units of $k_{B}T$ so the temperature will no longer enter our discussion. We simulate a collection of $N=nN_{c}$ monomers with interactions chosen such that $\delta_{ij}^{\vec{l}}>0$ if $i$ and $j$ are copies of neighbouring monomers with separation $\vec{l}$ in the target structure. We denote by $\rho$ the concentration of any specified monomer, which is assumed equal for all monomers. All our simulations are run in a box with side length $L$ and volume $V=L^{2}$, at volume fraction $\phi=N/V\approx 0.05$; note that $\phi=4^{k}\rho$.
The dynamics is modeled using the Virtual Move Monte Carlo algorithm Whitelam and Geissler (2007), whose equilibrium converges to the Boltzmann distribution, and which preserves many natural dynamical features when time is measured in Monte Carlo MC sweeps. Importantly, this algorithm allows moving clusters as a unit, merging clusters, and breaking apart clusters into sub-clusters. The average diffusion coefficient of a cluster may be a prescribed function of the cluster’s size or shape. We use an approximation to 3d Stokes’ drag, which implies a $4^{k}$-square diffuses approximately with diffusion coefficient $D_{k}\approx 1/\sqrt{n}=1/2^{k}$. To speed up the code, we enforce that monomers cannot rotate, which we expect not to change results qualitatively Evans and Winfree (2017).
@@
Figure 3: Building 4-squares with varying side energies $\delta_{1}$. Top left: the percentage of fully-completed 4-squares from simulations at different times (markers/solid lines). Black dashed line shows the equilibrium fraction of 4-squares, from Eqn. (12). Simulations considered $N_{c}=100$, $L=89$, and for each $\delta_{1}$ were repeated 12 times to obtain an average yield. Bottom: snapshots of individual simulations after 128 MC sweeps, showing about 1/9th of each system. For small $\delta_{1}$, the system is in equilibrium and contains a mixture of fragment sizes. For medium $\delta_{1}$, the system is nearly 100% 4-squares. For large $\delta_{1}$, the system forms a mixture of 3-mers and 4-squares. The 3-mers cannot fit together to form 4-squares without breaking apart, which happens on a longer timescale, so the system is kinetically trapped. Top right: the optimal interaction energy $\delta_{\rm opt}$ which maximizes $c_{4}(t)$, the concentration of 4-squares at time $t$, and the corresponding inverse fraction of errors $\Delta^{-1}_{\rm opt}(t)=(1-c_{4}(t)/\rho)^{-1}$, found by solving nonlinear rate equations (16) at $\phi=0.05$. Dashed lines are the best-fit lines on a log-linear and log-log scale respectively; their slopes are shown in the legends and are in good agreement with the unit slopes predicted by Eqns. (2), (3).
@@ II.2 Assembling 4-squares
Theory: An exact formulation of the thermodynamics and kinetics of the formation of 4-squares is presented in the SI. Here we emphasize key facts. In the regime of high yield, the dominant species at equilibrium is the fully formed 4-squares of concentration $c_{4}$, followed by individual monomers of concentration $c_{1}$. Thus the decrease in yield is $\Delta_{1}\approx c_{1}/\rho$ while $c_{4}\approx\rho$. Imposing that the reaction whereby 4 single monomers form a 4-square satisfies detailed balance leads to $c_{1}=e^{-\delta_{1}}c_{4}^{1/4}$, which implies:
@@
To test the time-dependence of these two scaling laws, we compute the value of $\delta_{1}$ which optimizes the concentration of 4-squares at different times $t$, by solving nonlinear rate equations for the different structures entering the assembly process (SI Secion B). The predicted scaling laws with $t$ are confirmed in Fig.3.
Figure 4: Building 16-squares with varying side energies $\delta_{1},\delta_{2}$. Colors show the percentage of fully-completed 16-squares after $128\times 10^{3}$ MC sweeps. The average yield is nearly 100% for a range of energies near $\delta_{1}=2\delta_{2}$ (white dotted line), while the yield is notably worse for identical interactions $\delta_{1}=\delta_{2}$ (black dotted line). Simulations considered $N_{c}=10$ copies of a 16-square, and for each value of $(\delta_{1},\delta_{2})$ were repeated 12 times to obtain an average yield.
@@ II.3 Assembling 16-squares is best done using hierarchical interactions
We explore whether this intuition is correct by varying the interaction energies $\delta_{1},\delta_{2}$ for a system of 16-squares (Figure 4). The highest average yield after a given fixed time is 97.5%, for $(\delta_{1},\delta_{2})=(8,4.5)$. The yield is not sensitive to these parameters: it is above 90% for a selection of parameters roughly in the range $(\delta_{1},\delta_{2})\in[6.5,8.5]\times[3.5,5]$. This is near the line $\delta_{2}=\delta_{1}/2$, validating our hypothesis that hierarchical interactions should lead to efficient assembly. It is notable that uniform interaction energies ($\delta_{1}{\approx}\delta_{2}$) never lead to such high yield over the time scales considered – the highest uniform yield is 82.5% for $\delta_{1}{=}\delta_{2}{=}6$, but most yields are much smaller.
Figure 5: Yield of perfect 256-squares in simulations of $N_{c}=4$ copies of a 256-square, initialized as a gas. Purple colours show the average yields at varying times as indicated in color using hierarchical energies ($\delta_{1}{=}2\delta_{2}{=}4\delta_{3}{=}8\delta_{4}$), and red colours show the average yields using uniform energies ($\delta_{1}{=}\delta_{2}{=}\delta_{3}{=}\delta_{4}$). Simulations for each $\delta_{1}$ were repeated 12 times for hierarchical interactions, and 10 times for uniform interactions.
Figure 6: Yields of $n{=}4^{k}$-squares using hierarchical interactions (markers/solid lines), as a function of strongest interaction energy $\delta_{1}$, calculated at times increasing by a factor of 8 for each generation (except for $k{=}5$ as described in the text). Dashed lines correspond to equilibrium approximation Eq.4.
Figure 7: Snapshots of individual simulations of assembling 256-squares at the best-assembling interaction energies: hierarchical ($\delta_{1}=8.75$, top) and uniform ($\delta_{1}=5.25$, bottom).
Figure 8: The mass distribution of fragments of varying size ($x$ axis) as a function of time ($y$ axis) as indicated in color, for the simulations described in Fig. 5. Parameters correspond to the best-assembling hierarchical energy ($\delta_{1}=8.75$, top), and the best-assembling uniform energy ($\delta_{1}=5.25$, bottom). For hierarchical interactions, most mass lies in fragments of sizes 4, 16, 64, and finally 256, with some mass also in multiples of these sizes. For uniform interactions, individual growth trajectories appear in the mass plot, consistent with the notion that clusters grow by adding monomers one by one. The color scheme is mass1/3 to highlight low-mass details.

### v2

@@ Abstract
There is evidence that the self-assembly of complex molecular systems often proceeds hierarchically, by first building subunits that later assemble in larger entities, in a process that can repeat multiple times. Yet, our understanding of this phenomenon and its performance is limited. Here we introduce a simple model for hierarchical addressable self-assembly, where interactions between particles can be optimised to maximise the fraction of a well-formed target structure, or yield. We find that a hierarchical strategy leads to an impressive yield up to at least five generations of the hierarchy, and does not require a cycle of temperatures as used in previous methods. High yield is obtained when the microscopic interaction decreases with the scale of units considered, such that the total interaction between intermediate structures remains identical at all scales. We provide thermodynamic and dynamical arguments constraining the interaction strengths where this strategy is effective. Overall, our work characterizes an alternative strategy for addressable self-assembly at a fixed temperature, and provides insight into the mechanisms sustaining hierarchical assembly in biological systems.
@@ I Introduction
The ribosome, the central machinery that translates messenger RNA into proteins, illustrates the surprising ability of multi-components systems to self-assemble. It is a protein complex made of $\sim 100$ proteins (in addition to ribosomal RNA), containing around $10^{4}$ amino acids. Considering the astronomical number of possible configurations these amino acids could take, how does this system self-assemble correctly? Clearly, this process is simplified by the fact that amino-acids first fold into proteins, which later form a complex. There is evidence that such a hierarchical strategy, where small units first form before assembling into larger ones, is at play in the folding of individual proteins itself Dill et al. (2008). In this view, proteins first nucleate secondary structures such as beta sheets or alpha helices Baldwin and Rose (1999) or more complex “foldons” Fersht (1997) which later organize into the full tertiary structure. Hierarchical self-assembly is also used in chemistry to obtain supra-molecular structures, typically held together with hydrogen bonds or Van der Waals interactions, much weaker than the covalent bonds organizing molecules Elemans et al. (2003). More generally, a hierarchical organization is central to a variety of complex systems and processes, from engineering to the composition of companies from smaller entities Simon (1962). Despite its importance and ubiquity, understanding the principles governing hierarchical assembly remains a challenge.
@@
Such principles could be used and tested in the context of “addressable” self-assembly, where each unit of a target structure is distinct and must assemble into a particular location Rothemund and Andersen (2012); Rothemund (2006); Sacanna et al. (2013); Wang et al. (2012); Rogers and Manoharan (2015); Rogers et al. (2016); Hong et al. (2017); Murugan et al. (2015a). Using DNA strands as building blocks with highly specific interactions, experimentalists have so realized highly complex structures containing up to tens of thousands of units Wei et al. (2012); Ke et al. (2012); Ong et al. (2017); Evans et al. (2024). A challenge is to find the particular building blocks King et al. (2024) or protocols Goodrich et al. (2021); Bupathy et al. (2022); Engel et al. (2023) such that the system assembles into a desired structure. Experiments operate in the regime where all pair interactions have roughly the same strength Hormoz and Brenner (2011), although the importance of displaying a range of interactions has been proposed in various contexts Jhaveri et al. (2024); Kaplan et al. (2014). When these interactions are strong, or equivalently the temperature is low, the system assembles many partially-formed fragments of the target structures, in a phenomenon known as “monomer starvation”. These fragments must break apart to assemble into copies of the target structure, but they cannot do so on accessible timescales – although the target structure has the lowest free energy, it is not kinetically accessible. In contrast, with weaker interactions the target state no longer has the lowest free energy. One solution, identified experimentally and later rationalized theoretically Reinhardt and Frenkel (2014); Jacobs et al. (2015) is to anneal temperature: start at a high temperature where nucleation is rare, then slowly lower temperature to grow the nuclei, repeating until one forms a target structure. This nucleation strategy can be enhanced if some parts of the assembly present a larger chemical potential Murugan et al. (2015b). While this protocol has worked well as a method to form one or a small number of copies of a target structure, it wastes monomers, since many monomers do not assemble into a target structure, and furthermore it requires a precise temperature protocol.
@@
Figure 1: Schematic of our model and the hierarchical assembly scheme. (a) Our model considers particles on a square lattice interacting with pairwise energies. Clusters can move as a unit and sometimes break apart. The example shown contains enough monomers to make $N_{c}{=}4$ copies of a 4-square, but only one 4-square has fully formed. (b) Building large squares with addressable complexity using hierarchical interactions. Square monomers with sticky sides assemble into 4-squares with strong edge interactions $\delta_{1}$, 4-squares assemble into 16-squares with weaker individual edge interactions $\delta_{2}$, 16-squares assemble into 64-squares using even weaker edge interactions $\delta_{3}$, and so on. We expect good assembly when the interactions are chosen hierarchically, as $\delta_{2}{=}\delta_{1}/2$, $\delta_{3}{=}\delta_{2}/2$, $\delta_{4}{=}\delta_{3}/2$, etc, for suitable values of the strongest energy $\delta_{1}$. This way, the energy for gluing together $4^{k}$-squares into $4^{k+1}$-squares ($\delta_{1}$), is the same as for gluing together monomers into 4-squares.
@@ II.1 Model
We consider a system of sticky squares (monomers) on a two-dimensional square lattice with periodic boundary conditions, as depicted in Fig. 1 (left). We assume that monomer $i$ interacts with monomer $j$ with an interaction energy 0 if the squares are not touching, $\infty$ if the squares overlap, and $\delta_{ij}^{\vec{l}}$ if the squares are exactly touching with separation $\vec{l}\in\{(1,0),(-1,0),(0,1),(0,-1)\}$. The dependence on $\vec{l}$ models directional interactions. The target structure is an $n$-square containing $n=2^{k}\times 2^{k}=4^{k}$ monomers (Figure 1) where $k$ is some integer, with each monomer in a distinct location. Our goal is to make $N_{c}$ copies of this target structure at a constant temperature $T$. We measure interaction energies in units of $k_{B}T$ so the temperature will no longer enter our discussion. We simulate a collection of $nN_{c}$ monomers with interactions chosen such that $\delta_{ij}^{\vec{l}}>0$ if $i$ and $j$ are copies of neighbouring monomers with separation $\vec{l}$ in the target structure. We denote by $\rho$ the concentration of any specified monomer, which is assumed equal for all monomers. All our simulations are run in a box with side length $L$ and volume $V=L^{2}$, at volume fraction $\phi=nN_{c}/V\approx 0.05$; note that $\phi=4^{k}\rho$.
The dynamics is modeled using the Virtual Move Monte Carlo algorithm Whitelam and Geissler (2007), whose equilibrium converges to the Boltzmann distribution, and which preserves many natural dynamical features when time is measured in Monte Carlo MC sweeps (one MC sweep is $nN_{c}$ Monte Carlo moves). Importantly, this algorithm allows moving clusters as a unit, merging clusters, and breaking apart clusters into sub-clusters. The average diffusion coefficient of a cluster may be a prescribed function of the cluster’s size or shape. We use an approximation to 3d Stokes’ drag, which implies a $4^{k}$-square diffuses approximately with diffusion coefficient $D_{k}\approx 1/\sqrt{n}=1/2^{k}$. (We verify in the SI, Section S4, that other models for diffusion do not alter the success of our proposed scheme.) To speed up the code, we enforce that monomers cannot rotate, which we expect not to change results qualitatively Evans and Winfree (2017). Consistent with the physics of micronscale particles, we treat the particles as “sticky” and do not include any barriers to reacting, so monomers interact as soon as they are in contact. We note however that including such reaction barriers gives another parameter for sculpting the kinetics of self-assembly Benoist and Sartori (2025).
Details of the algorithm and its implementation, including how diffusion coefficients are enforced, are described in Whitelam and Geissler (2007). Our code is adapted from Hedges and a version is available on Github at Holmes-Cerfon .
Our model bears similarities in its thermodynamic properties to other computational models used to study addressable self-assembly, e.g. Murugan et al. (2015a). It is notably different from most other studies in two aspects. One is its dynamics; many studies (with a handful of exceptions) only allow monomers to diffuse. A second is that we work in the canonical ensemble, with fixed particle number; rather than the grand canonical ensemble. The canonical ensemble is the setting that experiments operate in, and this choice allows us to address one of the key issues facing experimentalists, namely monomer starvation.
@@
Figure 2: Building 4-squares with varying side energies $\delta_{1}$. Top left: the percentage of fully-completed 4-squares from simulations at different times (markers/solid lines). Black dashed line shows the equilibrium fraction of 4-squares, from Eq.(S4). Simulations considered $N_{c}=100$, $L=89$, and for each $\delta_{1}$ were repeated 12 times to obtain an average yield. Bottom: snapshots of individual simulations after 128 MC sweeps, showing about 1/9th of each system. For small $\delta_{1}$, the system is in equilibrium and contains a mixture of fragment sizes. For medium $\delta_{1}$, the system is nearly 100% 4-squares. For large $\delta_{1}$, the system forms a mixture of 3-mers and 4-squares. The 3-mers cannot fit together to form 4-squares without breaking apart, which happens on a longer timescale, so the system is kinetically trapped. Top right: the optimal interaction energy $\delta_{\rm opt}$ which maximizes $c_{4}(t)$, the concentration of 4-squares at time $t$, and the corresponding inverse fraction of errors $\Delta^{-1}_{\rm opt}(t)=(1-c_{4}(t)/\rho)^{-1}$, found by solving nonlinear rate equations (S8) at $\phi=0.05$. Dashed lines are the best-fit lines on a log-linear and log-log scale respectively; their slopes are shown in the legends and are in good agreement with the unit slopes predicted by Eqs. (2), (3).
@@ II.2 Assembling 4-squares
Theory: An exact formulation of the thermodynamics and kinetics of the formation of 4-squares is presented in the SI, Section S1. Here we emphasize key facts. In the regime of high yield, the dominant species at equilibrium is the fully formed 4-squares of concentration $c_{4}$, followed by individual monomers of concentration $c_{1}$. Thus the decrease in yield is $\Delta_{1}\approx c_{1}/\rho$ while $c_{4}\approx\rho$. Imposing that the reaction whereby 4 single monomers form a 4-square satisfies detailed balance leads to $c_{1}=e^{-\delta_{1}}c_{4}^{1/4}$, which implies:
@@
To test the time-dependence of these two scaling laws, we compute the value of $\delta_{1}$ which optimizes the concentration of 4-squares at different times $t$, by solving nonlinear rate equations for the different structures entering the assembly process (SI Section S2). The predicted scaling laws with $t$ are confirmed in Fig.2.
Figure 3: Hierarchical assembly leads to high yield. (a) Building 16-squares with varying side energies $\delta_{1},\delta_{2}$. Colors show the percentage of fully-completed 16-squares after $128\times 10^{3}$ MC sweeps. The average yield is nearly 100% for a range of energies near $\delta_{1}=2\delta_{2}$ (white dotted line), while the yield is notably worse for identical interactions $\delta_{1}=\delta_{2}$ (black dotted line). Simulations considered $N_{c}{=}10$ copies of a 16-square, and for each value of $(\delta_{1},\delta_{2})$ were repeated 12 times to obtain an average yield. (b) Yield of perfect 256-squares in simulations of $N_{c}{=}4$ copies of a 256-square, for both hierarchical interactions ($\delta_{1}{=}2\delta_{2}{=}4\delta_{3}{=}8\delta_{4}$) and uniform interactions ($\delta_{1}{=}\delta_{2}{=}\delta_{3}{=}\delta_{4}$). The yield is plotted after $248\times 10^{5}$ MC sweeps, and simulations for each $\delta_{1}$ were repeated 12 times for hierarchical interactions, and 10 times for uniform interactions. (c) Yields of $n{=}4^{k}$-squares using hierarchical interactions (markers/solid lines), as a function of strongest interaction energy $\delta_{1}$, calculated at times increasing by a factor of 8 for each generation (except for k=5k{=}5 as described in the text). Dashed lines correspond to equilibrium approximation Eq.4. All simulations consider 2048 monomers, so the number of copies is $N_{c}=512,128,32,8,2$ for 4-, 16-, 64-, 256-, 1024-squares respectively. Systems of 64-, 256-, and 1024-squares are repeated four times and the yields are averaged. Vertical lines are one-standard deviation error bars.
@@ II.3 Assembling 16-squares is best done using hierarchical interactions
We explore whether this intuition is correct by varying the interaction energies $\delta_{1},\delta_{2}$ for a system of 16-squares (Figure 3 left). The highest average yield after a given fixed time is 97.5%, for $(\delta_{1},\delta_{2})=(8,4.5)$. The yield is not sensitive to these parameters: it is above 90% for a selection of parameters roughly in the range $(\delta_{1},\delta_{2})\in[6.5,8.5]\times[3.5,5]$. This is near the line $\delta_{2}=\delta_{1}/2$, validating our hypothesis that hierarchical interactions should lead to efficient assembly. It is notable that uniform interaction energies ($\delta_{1}{\approx}\delta_{2}$) never lead to such high yield over the time scales considered – the highest uniform yield is 82.5% for $\delta_{1}{=}\delta_{2}{=}6$, but most yields are much smaller.
Figure 4: Snapshots of individual simulations from Fig. 3 (b) of assembling 256-squares at the optimal interaction energies for each: hierarchical ($\delta_{1}=8.75$, top) and uniform ($\delta_{1}=5.25$, bottom). Each panel shows the full simulation domain.
Figure 5: The mass distribution of fragments of varying size ($x$ axis) as a function of time ($y$ axis) as indicated in color, for the simulations described in Fig. 3 (b). Parameters correspond to the best-assembling hierarchical energy ($\delta_{1}=8.75$, top), and the best-assembling uniform energy ($\delta_{1}=5.25$, bottom). For hierarchical interactions, most mass lies in fragments of sizes 4, 16, 64, and finally 256, with some mass also in multiples of these sizes. For uniform interactions, individual growth trajectories appear in the mass plot, consistent with the notion that clusters grow by adding monomers one by one. The color scheme is mass1/3 to highlight low-mass details.